# A matrix-free convex limiting framework for continuous Galerkin methods with nonlinear stabilization

### v1

@@ 5.3.1 Solid body rotation
The solid body rotation problem introduced by LeVeque [37] is widely used to assess the ability of numerical methods to capture smooth and discontinuous shapes of an exact solution to the linear advection equation
@@
with the velocity field ${\bf v}(x,y)=(0.5-y,x-0.5)^{\top}$. The initial and boundary conditions for this test are defined in [25, 37]. We stop computations at the final time $t=2\pi$, at which the exact solution coincides with the initial data. The invariant domain to be preserved is the interval $\mathcal{G}=[0,1]$.
@@
Figure 6: Numerical solutions to the solid body rotation problem [37] at $t=2\pi$ calculated using a uniform mesh, $\mathbb{Q}_{1}$ elements, $h=\frac{1}{128}$ and $\Delta t=10^{-3}$.
@@ 5.3.2 KPP problem
Extending our experiments to nonlinear scalar equations of the form (1a) in 2D, we consider the KPP problem [23], in which $\mathbf{f}(u)=(\sin(u),\cos(u))$ and $\Omega=(-2,2)\times(-2.5,1.5)$. The initial condition is given by
@@
At the final time $t=1$ the exact entropy solution of the KPP problem exhibits a spiral-shaped shock structure, which is captured correctly by WENO schemes only for the conservative choice $q=1$ of the steepening parameter (not shown here). To avoid convergence to spurious weak solutions for larger values of $q$, we replace the sensor $\gamma_{e}$ of formula (55) by $\min\{\gamma_{e},\xi_{e}\}$, where $\xi_{e}$ is an entropy correction factor. Using the nodal sensors $R_{i}$ of entropy viscosity stabilization employed in [9] and [34, Eq. 51], we set $\xi_{e}=\min_{i\in\mathcal{N}_{e}}R_{i}$.
@@
Figure 7: Numerical solutions to the KPP problem [23] at $t=1$ calculated using a uniform mesh, $\mathbb{Q}_{1}$ elements, $h=\frac{1}{128}$ and $\Delta t=10^{-3}$.
@@ 6 Conclusions
Even more significant efficiency gains can presumably be achieved for CG-LGL finite element schemes, because the nonuniform distribution of nodes within the cell makes the subcell CFL condition far more restrictive than that for the uniform submesh of Bernstein nodes [15]. Furthermore, the imposition of global bounds on LGL nodal states is less restrictive than the corresponding inequality constraints for Bernstein nodes. Hence, it is less likely that convex limiting will degrade the rates of convergence to smooth solutions. Adaptation of the proposed algorithms to the LGL setting is straightforward. Regardless of the basis, the performance of practical implementations can be greatly improved by exploiting the matrix-free nature of the presented methods and the data locality of high-order finite element approximations [4, 6, 22].

### v2

@@ 5.3.1 Solid body rotation
The solid body rotation problem introduced by LeVeque [leveque1996] is widely used to assess the ability of numerical methods to capture smooth and discontinuous shapes of an exact solution to the linear advection equation
@@
with the velocity field ${\bf v}(x,y)=(0.5-y,x-0.5)^{\top}$. The initial and boundary conditions for this test are defined in [kuzmin2020, leveque1996]. We stop computations at the final time $t=2\pi$, at which the exact solution coincides with the initial data. The invariant domain to be preserved is the interval $\mathcal{G}=[0,1]$.
@@
Figure 7: Numerical solutions to the solid body rotation problem [leveque1996] at $t=2\pi$ calculated using a uniform mesh, $\mathbb{Q}_{1}$ elements, $h=\frac{1}{128}$ and $\Delta t=10^{-3}$.
@@ 5.3.2 KPP problem
Extending our experiments to nonlinear scalar equations of the form (1a) in 2D, we consider the KPP problem [kurganov2007], in which $\mathbf{f}(u)=(\sin(u),\cos(u))$ and $\Omega=(-2,2)\times(-2.5,1.5)$. The initial condition is given by
@@
At the final time $t=1$ the exact entropy solution of the KPP problem exhibits a spiral-shaped shock structure, which is captured correctly by WENO schemes only for the conservative choice $q=1$ of the steepening parameter (not shown here). To avoid convergence to spurious weak solutions for larger values of $q$, we replace the sensor $\gamma_{e}$ of formula (58) by $\min\{\gamma_{e},\xi_{e}\}$, where $\xi_{e}$ is an entropy correction factor. Using the nodal sensors $R_{i}$ of entropy viscosity stabilization employed in [guermond2018] and [kuzmin2020a, Eq. 51], we set $\xi_{e}=\min_{i\in\mathcal{N}_{e}}R_{i}$.
@@
Figure 8: Numerical solutions to the KPP problem [kurganov2007] at $t=1$ calculated using a uniform mesh, $\mathbb{Q}_{1}$ elements, $h=\frac{1}{128}$ and $\Delta t=10^{-3}$.
5.4 Euler equations in 2D: double Mach reflection
Finally, we assess the performance of our scheme for multidimensional systems by running a 2D test for the compressible Euler equations. We restrict ourselves to computations with the WENO-L version using the full element contributions (45). In contrast to the two-dimensional scalar benchmarks of Sec. 5.3, we employ the default linear weights in the WENO reconstruction, namely $w_{l}^{e,\rm lin}=10^{-3}$ for $l=1,\ldots,n_{e}$ and $w_{0}^{e,\rm lin}=1-\sum_{l=1}^{n_{e}}w_{l}^{e,\rm lin}$. The steepening parameter of the smoothness sensor (7) is set to $q=1$.
As a benchmark problem, we consider the double Mach reflection test introduced by Woodward and Colella [woodward1984], which can be regarded as a two-dimensional extension of the blast-wave problem from Sec. 5.2.2. It represents a challenging test for multidimensional shock-capturing schemes because excessive diffusion smears the fine-scale features while insufficient stabilization leads to spurious oscillations. The computational domain is $\Omega=(0,4)\times(0,1)$. The boundary consists of a reflecting wall $\Gamma_{w}=\{(x,0)^{\top}\in\partial\Omega:1/6\leq x\leq 4\}$, a supersonic outlet $\Gamma_{\rm out}=\{(4,y)^{\top}:0<y\leq 1\}$, and a supersonic inlet $\Gamma_{\rm in}=\partial\Omega\setminus(\Gamma_{w}\cup\Gamma_{\rm out})$. The prescribed post-shock and pre-shock states are
$(\varrho_{L},v_{x,L},v_{y,L},p_{L})$ $=(8.0,8.25\cos(30^{\circ}),-8.25\sin(30^{\circ}),116.5),$
$(\varrho_{R},v_{x,R},v_{y,R},p_{R})$ $=(1.4,0.0,0.0,1.0),$
respectively. Initially, the post-shock state is imposed in $\Omega_{L}=\{(x,y)\;|\;x<\frac{1}{6}+\frac{y}{\sqrt{3}}\}$, and the pre-shock state in $\Omega_{R}=\Omega\setminus\Omega_{L}$. To account for the motion of the Mach $10$ shock along the top boundary, time-dependent inflow conditions are prescribed by imposing the post-shock state for $x<\frac{1}{6}+\frac{1+20t}{\sqrt{3}}$ and the pre-shock state elsewhere. The shock-wall interaction produces a Mach $10$ shock impinging on the reflecting wall at an angle of $60^{\circ}$.
In all simulations, the WENO-L method employs the pressure limiter from [kuzmin2023, kuzmin2020c], as described in Appendix A. The pressure fix proposed in [abgrall-arxiv, wissocq2025] produces virtually identical results (not shown here) in this test. No visible differences were observed between the two versions since the pressure correction is activated only in a very small number of cells.
The simulations are run up to the final time $t=0.2$. As DG methods are widely used for challenging hyperbolic problems, we compare our approach with the DG-WENO scheme from [vedral-arxiv] equipped with the classical Zhang–Shu slope limiter [zhang2010b]. For a fair comparison, both schemes use roughly the same number of degrees of freedom per variable, with only minor differences related to the boundary discretization.
The numerical results obtained with linear and quadratic finite elements are shown in Fig. 9. Remarkably, the quality of the WENO-L approximation is comparable to that of the DG-WENO solution, and the results remain free of spurious oscillations. This observation demonstrates the competitiveness of continuous finite element approaches for demanding compressible flow problems and may encourage further developments in this direction.
(a) WENO-L, $\mathbb{Q}_{1}$ elements, $N_{h}=591745$, $u_{h}=[1.238,22.415]$
(b) DG-WENO, $\mathbb{Q}_{1}$ elements, $N_{h}=589824$, $u_{h}=[1.400,22.116]$
(c) WENO-L, $\mathbb{Q}_{2}$ elements, $N_{h}=591745$, $u_{h}=[1.338,22.517]$
Figure 9: Density profiles for the double Mach reflection problem [woodward1984] at $t=0.2$ calculated using a uniform mesh.
@@ 6 Conclusions
Even more significant efficiency gains can presumably be achieved for CG-LGL finite element schemes, because the nonuniform distribution of nodes within the cell makes the subcell CFL condition far more restrictive than that for the uniform submesh of Bernstein nodes [hajduk2025a]. Furthermore, the imposition of global bounds on LGL nodal states is less restrictive than the corresponding inequality constraints for Bernstein nodes. Hence, it is less likely that convex limiting will degrade the rates of convergence to smooth solutions. Adaptation of the proposed algorithms to the LGL setting is straightforward. Regardless of the basis, the performance of practical implementations can be greatly improved by exploiting the matrix-free nature of the presented methods and the data locality of high-order finite element approximation [andrej2024, bello2020, kronbichler2019].
@@ Acknowledgments
Appendix A: IDP limiter for the Euler equations
Let $f=(f_{1},\ldots,f_{N})$ be an array of $N=2$ antidiffusive fluxes $(f_{ee^{\prime}},f_{e^{\prime}e})$ or $N=|\mathcal{N}_{e}|$ element contributions $(f_{j}^{e},\ j\in\mathcal{N}_{e})$. In both cases, the components of $f$ possess the zero sum property $\sum_{i=1}^{N}f_{i}=0$ and modify intermediate cell averages or nodal states $\bar{u}_{1},\ldots,\bar{u}_{N}\in\mathcal{G}$ as follows:
$\bar{u}_{i}^{*}=\bar{u}_{i}+\frac{\alpha f_{i}}{m_{i}}\in\mathcal{G},\qquad\alpha=\min_{1\leq i\leq N}\alpha_{i}.$
The scaling factors $m_{i}>0$ correspond to $|S_{ee^{\prime}}|$ in the flux limiter version and to $m_{i}^{e}$ in the slope limiter version. The choice of the upper bounds $\alpha_{i}\in[0,1]$ for $\alpha\in[0,1]$ should guarantee that $\bar{u}_{i}^{*}\in\mathcal{G}$ for $i=1,\ldots,N$. We emphasize that $i$ is not the global index of a node or element but the local index of an antidiffusive component $f_{i}$ to be limited using the correction factor $\alpha$.
For the Euler equations, $\bar{u}_{i}$ and $f_{i}$ are vectors with $d+2$ components:
$\bar{u}_{i}=(\rho_{i},(\rho\mathbf{v})_{i},(\rho E)_{i})^{\top},\qquad f_{i}=(f_{i}^{\rho},\mathbf{f}_{i}^{\rho\mathbf{v}},f_{i}^{\rho E})^{\top}.$
Positivity preservation for the density $\bar{\rho}_{i}^{*}$ can easily be enforced using
$\alpha_{i}^{\rho}=\begin{cases}-\frac{m_{i}\bar{\rho}_{i}}{f_{i}^{\rho}}&\mbox{if}\ m_{i}\bar{\rho}_{i}+f_{i}^{\rho}<0,\\ 1&\mbox{otherwise}.\end{cases}$
The pressure $p(\bar{u}_{i}^{*})$ and internal energy $e(\bar{u}_{i}^{*})$ are nonnegative if
$\overline{(\rho E)}_{i}^{*}\geq\frac{|\overline{(\rho\mathbf{v})}_{i}^{*}|^{2}}{2\bar{\rho}_{i}^{*}}.$
The corresponding inequality constraint for $\alpha$ can be written as (cf. [kuzmin2020, kuzmin2020c])
$P_{i}(\alpha)\geq Q_{i}=m_{i}\left(\frac{|\overline{(\rho\mathbf{v})}_{i}|^{2}}{2}-\bar{\rho}_{i}\overline{(\rho E)}_{i}\right),$
where
$P_{i}(\alpha)$ $=\alpha\left(\bar{\rho}_{i}f_{i}^{\rho E}+\overline{(\rho E)}_{i}f_{i}^{\rho}-\overline{(\rho\mathbf{v})}_{i}\cdot{\mathbf{f}}_{i}^{\rho\mathbf{v}}\right)+\frac{\alpha^{2}}{m_{i}}\left(f_{i}^{\rho E}f_{i}^{\rho}-\frac{|{\mathbf{f}}_{i}^{\rho\mathbf{v}}|^{2}}{2}\right).$
Note that $Q_{i}\leq 0$ because $\bar{u}_{i}\in\mathcal{G}$, where $\mathcal{G}\subseteq\{u\,:\,\rho(u)>0,\ e(u)>0\}$.
We can now derive a bound $\alpha_{i}^{p}$ for $\alpha$ following [kuzmin2020, kuzmin2020c] and [kuzmin2023, Examples 3.39, 4.16, 5.14]. For any $\alpha\in[0,1]$, we have $P_{i}(\alpha)\geq\alpha R_{i}$, where
$R_{i}$ $=\bar{\rho}_{i}f_{i}^{\rho E}+\overline{(\rho E)}_{i}f_{i}^{\rho}-\overline{(\rho\mathbf{v})}_{i}\cdot{\mathbf{f}}_{i}^{\rho\mathbf{v}}+\frac{1}{m_{i}}\min\Big\{0,f_{i}^{\rho E}f_{i}^{\rho}-\frac{|{\mathbf{f}}_{i}^{\rho\mathbf{v}}|^{2}}{2}\Big\}.$
To ensure that $\alpha R_{i}\geq Q_{i}$ for $i\in\{1,\ldots,N\}$, the pressure limiter must use
$\alpha_{i}^{p}=\begin{cases}\frac{Q_{i}}{R_{i}}&\mbox{if}\ R_{i}<Q_{i},\\ 1&\mbox{otherwise}.\end{cases}$
We remark that this formula is very similar to the one derived in [abgrall-arxiv, wissocq2025] using the general framework of geometric quasi-linearization [wu2023]. Both approaches replace the nonlinear pressure constraint by linear sufficient conditions.